If you were to randomly select 3 movies from a selection of 29 movies, how many different lists of 3 could you have?
Re-arranging the 3 does NOT make a new list.
I tried to do 29*28*27 which = 21924 but I think that that includes all the 3 groups that are re-arranged.
How should I go about solving this problem?

Answer by stanbon(75887)

If you were to randomly select 3 movies from a selection of 29 movies, how many different lists of 3 could you have?
------------------
You were counting arrangements of three objects.
You want to count GROUPS of three objects.
29C3 = 29!/[(29-3)!*3!]
= (29*28*27)/(1*2*3) = 29*14*9 = 3654
=========================
Cheers,
Stan H.